# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v3

@@ 1 Introduction
Understanding the behavioral effects of reference points, particularly in the context of pressure and risk-taking, is a critical area of research. The core concept is that an individual’s assessment of an outcome is influenced not only by the outcome itself but also by how it compares to a reference point (Tversky and Kahneman 1992). A significant challenge in this field is disentangling the effects of pressure from a reference point on risk-taking behavior and outcomes. Unlike controlled experimental settings as in Schwerter (2024), where risk-taking can be modeled as a binary lottery, field data presents complexities that make such decompositions difficult. In this paper, we utilize a sequential competition setting and panel dataset from official bench press competitions, offering an opportunity to explore the interplay between pressure and risk-taking behavior. This dataset enables us to observe weight attempts, outcomes, rankings, and the exogenous pressure exerted by rivals over time. By examining how pressure influences subsequent risk-taking decisions, we aim to enhance the understanding of behavior under uncertainty and contribute to discussions on optimizing competition design to maximize performance.
The primary contribution of this study is the decomposition of the effects of pressure arising from reference points into distinct elements: the choice of weight attempt (analogous to a lottery choice) and the probability of success (analogous to a lottery gain). We also explore alternative competition designs that manipulate the availability of pressure-related information. The structure of official bench press competitions provides an ideal context to study the relationship between reference points and risk-taking within the sequential game setting of weight attempts and weight lifting stages, where exogenous pressure is present. Specifically, a weight attempt under pressure from rivals can be seen as a risk-taking behavior similar to making a lottery choice. Whether the lift is successful depends on the pressure conditional on the weight attempt and the rivals’ outcomes. Through this investigation, we seek to answer an analogous question in real-world competitive settings: What would be the impact on risk-taking behavior and expected outcomes if the pressure exerted by rivals were removed?
Our results demonstrate the role of pressure in shaping lifters’ behavior during competition. Pressure from both lower and higher-ranked rivals leads to increased attempt weights across both the second and third attempts, reflecting lifters’ tendency to challenge themselves under competition.
@@
Heterogeneity analyses further reveal that responses to pressure vary by gender, experience, and rivalry history. Male lifters exhibit stronger responses to both lower and higher rival pressure in attempt selection, whereas female lifters show a more muted response, especially under upward pressure. Female lifters experience a sharper decline in success probability when attempting to turn around a higher-ranked rival, suggesting a heightened sensitivity to stress. More experienced lifters are more sensitive to higher-ranked competitors, reinforcing the idea that competitive awareness strengthens with experience. Additionally, lifters with a history of repeated rivalry encounters show heightened sensitivity to pressure, suggesting that familiarity enhances their ability to navigate competition. For success probabilities, while pressure from lower-ranked rivals can improve outcomes, particularly among experienced lifters, upward pressure generally reduces them, with stronger effects among female competitors. Lifters with strong rivalry history show a small negative impact from upward pressure but a large positive impact from lower pressure. These confirm substantial heterogeneity.
To examine how pressure shapes the balance between ambition and feasibility in decision-making, we simulate counterfactual competition designs in which such pressure is removed. The results reveal consistent directional patterns—more conservative attempts—but also considerable heterogeneity in expected outcomes. While a subset of lifters benefit from the absence of pressure, some lifters experience lower success rates and reduced expected outcomes. These differences reflect the uneven role of external pressure: for some, it enables focus and composure; for others, it disrupts execution or leads to overly cautious choices.
@@ 1.1 Related literature
First, our paper contributes to the growing body of literature on reference points in sports competition. We follow the framework of O’Donoghue and Sprenger (2018), who suggest a variety of reference points, including the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other possibilities such as norms and social comparisons. However, the determination of reference points is often arbitrary and context-dependent (DellaVigna 2018), with these effects being composited and mixed in actual sports settings. [^1] Baillon et al. (2020) use Bayesian hierarchical modeling to estimate the marginal posterior distributions of six reference point rules. They found that subjects most frequently used the status quo and MaxMin as reference points. Table 1 summarizes the data, reference points, classifications, and the loss-gain domain of the utility function in related studies, although the classification is not perfect.
@@
Salient numbers, such as round finish times in marathons (Allen et al. 2017, Soetevent 2022) and predetermined par scores in golf (Pope and Schweitzer 2011), are well-known as focal reference points. Runners adjust their effort levels to meet target times, and golf players are significantly less accurate when attempting shots ahead of a predetermined reference point, labeled as “par.”  [^2] Beyond sports competition, Rees-Jones (2018) provides evidence implying that individuals have reference-dependent preferences, with zero tax due serving as the reference point. As experimental evidence, Abeler et al. (2011) conducted an experiment manipulating the fixed pay amount as a reference point in a real-effort task. In this experiment, either fixed pay or piecework pay for a correct answer was randomly determined after task completion. They found that subjects were more likely to stop when the expected piecework pay equaled the fixed pay, exerting more effort when the fixed pay was high. Corgnet et al. (2015) also show that goal setting by managers is most effective for team workers when monetary incentives are strong. Goals set by athletes based on past or current performances are treated as reference points in the form of status quo, past experience, and expectations. For example, self-reported target times in marathons (Markle et al. 2018) and personal bests in chess (Anderson and Green 2018, González-Díaz et al. 2024) serve as natural reference points, whether or not they function explicitly as goals. Theoretically, Heath et al. (1999) and Koch and Nafziger (2016) discuss the concept of goals as reference points. In our data on bench press competition, we do not observe prevalent bunching around round weights, likely because the categories in which players participate are conditional on characteristics such as age, body weight, and gender. Additionally, unlike professional sports that provide substantial rewards, bench press competitions are amateur events with no direct monetary incentives. This setting highlights lifters’ self-motivation to surpass personal bests and rank updates based on recent outcomes, independent of financial incentives. The current rank is the reference point and each lifter has more incentive to avoid a large loss from decreasing his ranking rather than to get a relatively small gain from increasing his ranking.
@@
In terms of pressure from rivals’ outcomes, the phenomenon of “choking under pressure”—where performance declines under stress—has been extensively studied. For instance, large-sample studies reject the hypothesis that first-mover advantage in soccer or being slightly behind increases the likelihood of winning in basketball, football, or rugby (Kocher et al. 2012, Teeselink et al. 2023). [^3] Similar studies include the effects of spectators on penalty shootouts in soccer (Dohmen 2008), free throws and playoff performance in basketball (Hickman and Metz 2015, Böheim et al. 2019, Cao et al. 2011, Toma 2017, Morgulev and Galily 2018), the impact of bookmakers’ odds on soccer play (Bartling et al. 2015), putting performance in golf (Hickman and Metz 2015), the influence of decisive moments on dart throws (Teeselink et al. 2020), and biathlon performance (Harb-Wu and Krumer 2019, Lindner 2017), the effect of previous game outcomes on current outcomes in hockey (Kniffin and Mihalek 2014), and gender differences in pressure in tennis (Cohen-Zada et al. 2017, Paserman 2023). While some papers implicitly link the cause of pressure to reference points, we do not explicitly classify pressure under the loss-gain utility framework, as the specific cases do not fit neatly within that domain. Our study also relates to peer effects in terms of externalities from rivals’ abilities and performance. Guryan et al. (2009), using random assignment in high-stakes golf tournaments, find no evidence of peer effects on average. In contrast, Yamane and Hayashi (2015) observe positive peer effects in swimming competitions, particularly suggesting that being chased by peers with lower personal bests enhances performance. These mixed results indicate that peer effects vary across different contexts, with Nishihata (2022) showing that peer effects among speed skaters depend on race distance and prize money. In our study, we leverage the sequential timing of observing rivals’ outcomes, applying the attempt weight, and executing the lift to decompose the effects of rivals’ outcomes on both the weight selection and the success probability of the lift.
@@
Second, this paper contributes to the literature on the effect of reference points on risk-taking actions. The most closely related studies are Puente-Moncayo (2019) and Genakos and Pagliero (2012). Puente-Moncayo (2019) examines the effect of current scores on the tendency to take high-risk shots in tennis. The author follows Ely et al. (2017), using the approach developed by Klaassen and Magnus (2009) (a structural model) to show that first serves should be riskier than second serves, and using two first serves (or two second serves) is suboptimal. Genakos and Pagliero (2012) study the effects of interim ranking in itself on weight application and success probability using the data of the powerlifting competition in the Olympic Games and World and European Championships between 1990 and 2006. In the experimental economics literature, Schwerter (2024) conducted experiments on risk-taking behavior when subjects were exposed to exogenously predetermined peer earnings. The author found that subjects exposed to higher peer earnings increased their risk-taking behavior. Similarly, Post et al. (2008) highlight the relevance of risk-taking behavior in path-dependent contexts, using data from the TV show “Deal or No Deal” and a replicated experiment where subjects choose to either “deal” for a sure prize or “no deal” to continue for an uncertain prize. Our paper is more ideal than these examples for several reasons. First, the weight selection conditional on personal bests and rivals’ outcomes is analogous to choosing a lottery, where the probability of successfully lifting the weight can be inferred from data by both lifters and econometricians. Second, lifting weights involves a simple, monoarticular action, which leads to relatively fewer external factors influencing the outcomes, compared to other sports including powerlifting (Genakos and Pagliero 2012). Third, unlike Genakos and Pagliero (2012), we exploit the sequential competition design which explicitly defines revealed information to each player. Fourth, we use a universe of official bench press competitions with a richer sample size and set of variables, which have been known recently as the most important factors to detect behavioral effects in the sports data (Teeselink et al. 2023).
@@
Third, our paper contributes to the literature on competition design aimed at maximizing each player’s score within a group. This contribution parallels general frameworks in competition design within organizations and is related to contest theory, although sabotage and uncooperative behavior, which can occur in real-world workplace competition, are less likely in sports competition. [^4] Drago and Garvey (1998) find that strong promotion incentives reduce helping behavior, even though individual effort increases. Bandiera et al. (2005) show that fruit pickers reduce their effort under relative incentives when their peers are friends, due to the negative externality of individual effort on others’ payoffs. As outlined by Lazear and Oyer (2013), relative performance evaluation (RPE) exists in two forms. The first allows firms to filter out common shocks by comparing an individual’s performance to that of a peer group. Gibbons and Murphy (1990) provide empirical evidence suggesting that RPE was used in the compensation contracts of U.S. CEOs from 1974 to 1986. The second form of RPE involves fixed rewards based on the performance rankings of participants, commonly referred to as a rank-order tournament. In tournament theory, as modeled by Lazear and Rosen (1981), large differences in payoffs across ranks motivate those at lower levels to exert more effort, while some risk-averse agents may exert less effort when the importance of winning is diminished. Rank-order tournaments have a wide range of applications, including sports competitions. For example, consistent with tournament theory, Brown (2011) shows that the presence of a superstar highly likely to win reduces the performance of other players. In our context, the unique structure of official bench press competitions allows us to compare alternative competition settings to the current format through simulations based on our estimates.
@@
Paper Data Reference Point Classification Loss-Gain Domain
Pope and Schweitzer (2011) Golf Par For Each Hole Focal Outcomes Score For Each Hole
Pope and Simonsohn (2011) Basketball Round Numbers In Performance Focal Outcomes Current Outcome
Allen et al. (2017) Marathon Round Finish Time Focal Outcomes Finish Time
Essl and Jaussi (2017) Lab Experiment Points (ECU) When <,><,>< , > 10sec Focal Outcomes Points From Real-Effort Task
Soetevent (2022) Running Event Round Finishing Time Focal Outcomes Finishing Time
Schwerter (2024) Lab Experiment Peer Earnings Focal Outcomes Earnings From Binary Lotteries
Post et al. (2008) TV Shows Earlier Expectations Expectations Bank Offer
Abeler et al. (2011) Lab Experiment Amount Of Fixed Pay Expectations Piece Rate For Correct Answers
Bartling et al. (2015) Soccer Expected Outcome By Odds Expectations Current Outcome
Markle et al. (2018) Marathon Self-Reported Target Time Status Quo Finish Time
Kniffin and Mihalek (2014) Hockey The Outcomes For Game 1 Past Experience Score In Game 2
Anderson and Green (2018) Chess Online Personal Best Rating Past Experience Rating
González-Díaz et al. (2024) Chess In-Person Personal Best Rating Past Experience Rating
Berger and Pope (2011) Basketball Performance Of The Opponent Social Comparison Current Outcome
Puente-Moncayo (2019) Tennis Rival’s Score Social Comparison Score
Teeselink et al. (2023) Four Sports Performance Of The Opponent Social Comparison Current Outcome
This Paper Bench Press Current Own Ranking Social Comparison Ranking
Baillon et al. (2020) Lab Experiment 6 Rules Mixed Lottery Options
Notes: We follow the referent list of O’Donoghue and Sprenger (2018) which suggest a number of possibilities: the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other candidates including norms and social comparisons, although the reference point determination is, in ways, arbitrary and context-dependent (DellaVigna 2018) and these effects are composited and mixed in the actual sports.

### v4

# arXiv:2409.13333v4[econ.GN] 02 Mar 2026

@@ 1 Introduction
Understanding the behavioral effects of reference points, particularly in the context of pressure and risk-taking, is a critical area of research. The core concept is that an individual’s assessment of an outcome is influenced not only by the outcome itself but also by how it compares to a reference point (Tversky and Kahneman 1992). A significant challenge in this field is disentangling the effects of pressure associated with reference points on risk-taking behavior and outcomes. Unlike controlled experimental settings as in Schwerter (2024), where risk-taking can be modeled as a binary lottery, field data present complexities that make such decompositions difficult. In this paper, we utilize a sequential competition setting and panel dataset from official bench press competitions to explore the interplay between competitive pressure arising from reference points and risk-taking behavior. This dataset enables us to observe weight attempts, outcomes, rankings, and the exogenous pressure exerted by rivals over time. In particular, we treat a lifter’s interim ranking during the competition as the relevant reference point, which naturally emerges from the sequential structure of attempts. This implies that potential changes in rank are experienced as gains or losses, thereby generating competitive pressure. By examining how pressure influences subsequent risk-taking decisions, we aim to enhance the understanding of behavior under uncertainty and contribute to discussions on optimizing competition design to maximize performance.
The primary contribution of this study is the decomposition of the effects of pressure arising from reference points into distinct elements: the choice of weight attempt (analogous to a lottery choice) and the probability of success (analogous to a lottery gain). In this setting, choosing a heavier weight corresponds to selecting a lottery with a lower probability of success but a higher potential payoff in terms of achieved ranking, whereas choosing a lighter weight implies the opposite trade-off. We therefore interpret risk-taking as the selection of heavier attempt weights conditional on a lifter’s ability. Against this backdrop, the structure of official bench press competitions provides an ideal context to study the relationship between reference points and risk-taking within the sequential game setting of weight declaration and lifting stages, where exogenous pressure is present. Whether the lift is successful depends not only on the chosen weight but also on the competitive pressure arising from the possibility that the lifter’s relative position may change during the competition. We also explore alternative competition designs that manipulate the availability of pressure-related information. Through this investigation, we seek to answer an analogous question in real-world competitive settings: What would be the impact on risk-taking behavior and expected outcomes if the pressure exerted by rivals were removed?
Our results demonstrate the role of pressure in shaping lifters’ behavior during competition. Pressure from both lower- and higher-ranked rivals leads to increased attempt weights across both the second and third attempts, reflecting lifters’ tendency to challenge themselves under competition.
@@
Heterogeneity analyses further reveal that responses to pressure vary by gender, experience, and rivalry history. Male lifters exhibit stronger responses to pressure from both lower- and higher-ranked rivals in attempt selection, whereas female lifters show a more muted response, especially under upward pressure. Female lifters experience a sharper decline in success probability when attempting to overtake a higher-ranked rival, suggesting a heightened sensitivity to stress. More experienced lifters are more sensitive to higher-ranked competitors, reinforcing the idea that competitive awareness strengthens with experience. Additionally, lifters with a history of repeated rivalry encounters show heightened sensitivity to pressure, suggesting that familiarity enhances their ability to navigate competition. For success probabilities, while pressure from lower-ranked rivals can improve outcomes, particularly among experienced lifters, upward pressure generally reduces them, with stronger effects among female competitors. Lifters with strong rivalry history show a small negative impact from upward pressure but a large positive impact from lower pressure. These findings confirm substantial heterogeneity.
To examine how pressure shapes the balance between ambition and feasibility in decision-making, we simulate counterfactual competition designs in which such pressure is removed. The results reveal consistent directional patterns—more conservative attempts—but also considerable heterogeneity in expected outcomes. While a subset of lifters benefit from the absence of pressure, others experience lower success rates and reduced expected outcomes. These differences reflect the uneven role of external pressure: for some, it enables focus and composure; for others, it disrupts execution or leads to overly cautious choices.
@@ 1.1 Related literature
First, our paper contributes to the growing body of literature on reference points in sports competition. We follow the framework of O’Donoghue and Sprenger (2018), who suggest a variety of reference points, including the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other possibilities such as norms and social comparisons. However, the determination of reference points is often arbitrary and context-dependent (DellaVigna 2018), with these effects often combined in actual sports settings. [^1] Baillon et al. (2020) use Bayesian hierarchical modeling to estimate the marginal posterior distributions of six reference point rules. They found that subjects most frequently used the status quo and MaxMin as reference points. Table 1 summarizes the data, reference points, classifications, and the loss-gain domain of the utility function in related studies, although the classification is not perfect.
@@
Salient numbers, such as round finish times in marathons (Allen et al. 2017, Soetevent 2022) and predetermined par scores in golf (Pope and Schweitzer 2011), are well-known as focal reference points. Runners adjust their effort levels to meet target times, and golfers are significantly less accurate when attempting shots ahead of a predetermined reference point, labeled as “par.”  [^2] Beyond sports competition, Rees-Jones (2018) provides evidence implying that individuals have reference-dependent preferences, with zero tax due serving as the reference point. As experimental evidence, Abeler et al. (2011) conducted an experiment manipulating the fixed pay amount as a reference point in a real-effort task. In this experiment, either fixed pay or piecework pay for a correct answer was randomly determined after task completion. They found that subjects were more likely to stop when the expected piecework pay equaled the fixed pay, exerting more effort when the fixed pay was high. Corgnet et al. (2015) also show that goal setting by managers is most effective for team workers when monetary incentives are strong. Goals set by athletes based on past or current performances are treated as reference points in the form of status quo, past experience, and expectations. For example, self-reported target times in marathons (Markle et al. 2018) and personal bests in chess (Anderson and Green 2018, González-Díaz et al. 2024) serve as natural reference points, whether or not they function explicitly as goals. Theoretically, Heath et al. (1999) and Koch and Nafziger (2016) discuss the concept of goals as reference points. In our data on bench press competition, we do not observe prevalent bunching around round weights, likely because the categories in which lifters compete are narrowly defined by characteristics such as age, body weight, and gender, so that competition is localized within relatively small groups and ranking differences are often determined by small weight margins. Additionally, unlike professional sports that provide substantial rewards, bench press competitions are amateur events with no direct monetary incentives. This setting highlights lifters’ self-motivation to surpass personal bests and rank updates based on recent outcomes, independent of financial incentives. The current rank is the reference point, and each lifter is expected to have stronger incentives to avoid a large loss from a rank decrease than to pursue a relatively small gain from a rank increase.
@@
In terms of pressure from rivals’ outcomes, the phenomenon of “choking under pressure”—where performance declines under stress—has been extensively studied. For instance, large-sample studies reject the hypothesis that first-mover advantage in soccer or being slightly behind increases the likelihood of winning in basketball, football, or rugby (Kocher et al. 2012, Teeselink et al. 2023). [^3] Similar studies include the effects of spectators on penalty shootouts in soccer (Dohmen 2008), free throws and playoff performance in basketball (Hickman and Metz 2015, Böheim et al. 2019, Cao et al. 2011, Toma 2017, Morgulev and Galily 2018), the impact of bookmakers’ odds on soccer play (Bartling et al. 2015), putting performance in golf (Hickman and Metz 2015), the influence of decisive moments on dart throws (Teeselink et al. 2020), and biathlon performance (Harb-Wu and Krumer 2019, Lindner 2017), the effect of previous game outcomes on current outcomes in hockey (Kniffin and Mihalek 2014), and gender differences in pressure in tennis (Cohen-Zada et al. 2017, Paserman 2023). While some papers implicitly link the cause of pressure to reference points, we do not explicitly classify pressure under the loss-gain utility framework, as the specific cases do not fit neatly within that domain. Our study also relates to peer effects in terms of externalities from rivals’ abilities and performance. Guryan et al. (2009), using random assignment in high-stakes golf tournaments, find no evidence of peer effects on average. In contrast, Yamane and Hayashi (2015) observe positive peer effects in swimming competitions, particularly suggesting that being chased by peers with lower personal bests enhances performance. These mixed results indicate that peer effects vary across different contexts, with Nishihata (2022) showing that peer effects among speed skaters depend on race distance and prize money. In our study, we leverage the sequential timing of observing rivals’ declarations and outcomes, declaring the attempt weight, and executing the lift to decompose the effects of rivals’ declarations and outcomes on both weight selection and the success probability of the lift.
@@
Second, this paper contributes to the literature on the effect of reference points on risk-taking actions. The most closely related studies are Puente-Moncayo (2019) and Genakos and Pagliero (2012). Puente-Moncayo (2019) examines the effect of current scores on the tendency to take high-risk shots in tennis. The author follows Ely et al. (2017), using the approach developed by Klaassen and Magnus (2009) (a structural model) to show that first serves should be riskier than second serves, and using two first serves (or two second serves) is suboptimal. Genakos and Pagliero (2012) study the effects of interim ranking itself on weight selection and success probability using data from powerlifting competitions in the Olympic Games and World and European Championships between 1990 and 2006. In the experimental economics literature, Schwerter (2024) conducted experiments on risk-taking behavior when subjects were exposed to exogenously predetermined peer earnings. The author found that subjects exposed to higher peer earnings increased their risk-taking behavior. Similarly, Post et al. (2008) highlight the relevance of risk-taking behavior in path-dependent contexts, using data from the TV show “Deal or No Deal” and a replicated experiment where subjects choose to either “deal” for a sure prize or “no deal” to continue for an uncertain prize. Our setting is better suited than these examples for several reasons. First, weight selection conditional on personal bests and rivals’ outcomes is analogous to choosing a lottery, where the probability of successfully lifting the weight can be inferred from data by both lifters and econometricians. Second, lifting weights involves a simple, monoarticular action, which leads to relatively fewer external factors influencing outcomes compared with other sports, including powerlifting (Genakos and Pagliero 2012). Third, unlike Genakos and Pagliero (2012), we exploit the sequential competition design, which explicitly defines the information revealed to each lifter. Fourth, we use a universe of official bench press competitions with a richer sample size and set of variables, which have recently been recognized as important for detecting behavioral effects in sports data (Teeselink et al. 2023).
@@
Third, our paper contributes to the literature on competition design aimed at maximizing each lifter’s score within a group. This contribution parallels general frameworks in competition design within organizations and is related to contest theory, although sabotage and uncooperative behavior, which can occur in real-world workplace competition, are less likely in sports competition. [^4] Drago and Garvey (1998) find that strong promotion incentives reduce helping behavior, even though individual effort increases. Bandiera et al. (2005) show that fruit pickers reduce their effort under relative incentives when their peers are friends, due to the negative externality of individual effort on others’ payoffs. As outlined by Lazear and Oyer (2013), relative performance evaluation (RPE) exists in two forms. The first allows firms to filter out common shocks by comparing an individual’s performance to that of a peer group. Gibbons and Murphy (1990) provide empirical evidence suggesting that RPE was used in the compensation contracts of U.S. CEOs from 1974 to 1986. The second form of RPE involves fixed rewards based on the performance rankings of participants, commonly referred to as a rank-order tournament. In tournament theory, as modeled by Lazear and Rosen (1981), large differences in payoffs across ranks motivate those at lower levels to exert more effort, while some risk-averse agents may exert less effort when the importance of winning is diminished. Rank-order tournaments have a wide range of applications, including sports competitions. For example, consistent with tournament theory, Brown (2011) shows that the presence of a superstar highly likely to win reduces the performance of other participants. In our context, the unique structure of official bench press competitions allows us to compare alternative competition settings to the current format through simulations based on our estimates.
@@
Paper Data Reference Point Classification Loss-Gain Domain Pope and Schweitzer (2011) Golf Par For Each Hole Focal Outcomes Score For Each Hole Pope and Simonsohn (2011) Basketball Round Numbers In Performance Focal Outcomes Current Outcome Allen et al. (2017) Marathon Round Finish Time Focal Outcomes Finish Time Essl and Jaussi (2017) Lab Experiment Points (ECU) When <,><,> 10sec Focal Outcomes Points From Real-Effort Task Soetevent (2022) Running Event Round Finishing Time Focal Outcomes Finishing Time Schwerter (2024) Lab Experiment Peer Earnings Focal Outcomes Earnings From Binary Lotteries Post et al. (2008) TV Shows Earlier Expectations Expectations Bank Offer Abeler et al. (2011) Lab Experiment Amount Of Fixed Pay Expectations Piece Rate For Correct Answers Bartling et al. (2015) Soccer Expected Outcome By Odds Expectations Current Outcome Markle et al. (2018) Marathon Self-Reported Target Time Status Quo Finish Time Kniffin and Mihalek (2014) Hockey The Outcomes For Game 1 Past Experience Score In Game 2 Anderson and Green (2018) Chess Online Personal Best Rating Past Experience Rating González-Díaz et al. (2024) Chess In-Person Personal Best Rating Past Experience Rating Berger and Pope (2011) Basketball Performance Of The Opponent Social Comparison Current Outcome Puente-Moncayo (2019) Tennis Rival’s Score Social Comparison Score Teeselink et al. (2023) Four Sports Performance Of The Opponent Social Comparison Current Outcome This Paper Bench Press Current Own Ranking Social Comparison Ranking Baillon et al. (2020) Lab Experiment 6 Rules Mixed Lottery Options
Notes: We follow the referent list of O’Donoghue and Sprenger (2018), which suggests a number of possibilities: the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other candidates including norms and social comparisons, although reference point determination is in many ways arbitrary and context-dependent (DellaVigna 2018), and these effects are often combined in actual sports settings.